OpenStudy (anonymous):

Stacy made a deposit of $8,500.00 into an account. After ten years, the account had earned $5,525.00 in interest. What simple interest rate did the account earn? 2 4.5 5 6.5

OpenStudy (anonymous):

SI=PTR/100 Amount P=8500 term T=10years rate r=? 8500+5525=8500*10*r/100 14025=8500*10*r/100 r=14025*100/(8500*10) rate r=16.5 Answers in choice is non of these